# Unchosen Experts Can Contribute Too: Unleashing MoE Models’ Power by Self-Contrast

### v1

@@ 1 Introduction
Scaling up model parameters, dataset size and training time has been considered the most direct and effective approach to improving foundation models’ performance [18; 26; 29]. However, scaling dense models substantially increases computational costs, which poses a significant practical challenge. Mixture-of-Experts (MoE) [30; 39; 17; 9; 37; 34] has emerged as a compelling solution for optimizing the balance between model capacity and computation overhead in the era of large foundation models.
MoE models achieve the goal by sparsely activating only a portion of the parameters for each specific input. Specifically, in MoE models, parameters are grouped into a bunch of experts, MoE models only activate some of them for processing a given input. This selective activation is achieved through a routing mechanism that dispatches each input token to a fixed number of experts (e.g, top-$k$ routing [12; 10; 37; 17]). Therefore, compared to their dense counterparts, MoE models enjoy more efficient training with significantly reduced computational costs [10; 39; 17; 9; 37; 34]. At the inference stage, they typically adhere to the same routing strategy as the training stage, activating only a small fraction of experts. Basically, for each input token, most of the well-trained experts do not contribute to the output prediction. As a result, the potential of utilizing more experts during the inference stage to enhance performance remains underexplored.
@@
In this paper, we investigate the impact of unchosen experts111 Unchosen experts refer to the experts not selected by default routing (e.g., top-2 routing in Mixtral 8x7B). on the performance of MoE models and explore their suitable usage. A direct hypothesis is that incorporating more experts improves MoE models and helps solve more difficult problems [3; 36; 14]. However, in our exploratory experiment on Mixtral 8x7B [17], we find simply raising the number of activated experts (blue lines in Figure 1) does not lead to stable improvements and may even hurt performance on different tasks. This indicates that unchosen experts may contribute little or even negatively to the final performance, which is contrary to the common perception of unchosen experts as candidates of positive power.
@@ 2.2 Divergence Between Different Routing Strategies: An Exploratory Analysis
To sum up, it is observed that, in scenarios demanding reasoning capability for next-token prediction, MoE models with top-2 and rank-$k$ routing strategies demonstrate distinct generation behaviors. This phenomenon suggests that different experts do not always act synergistically, and could in fact exhibit conflicting behaviors. To harness such information introduced by more experts, a feasible approach is to apply contrastive methods [22; 20] to transform the observed negative impacts into positive ones.
@@ 3.3 Results
SCMoE enhances mathematical reasoning by a +5.10 increase on GSM8K, commonsense reasoning by a +3.46 improvement on StrategyQA. Moreover, in code generation, SCMoE gets improvements of +2.60 and +7.92 on the MBPP and HumanEval, respectively. In contrast, traditional search-based methods do not demonstrate substantial improvements on MoE models. In particular, DoLa’s performance not only fails to surpass, but actually falls below the greedy baseline, particularly due to its inability to terminate generation sequences appropriately (for specific examples, refer to Table 11 in the appendix). Meanwhile, contrastive decoding with Mistral 7B as the amateur model does not result in consistent improvements, and even a decrease in pass@1 accuracy on MBPP (-1.00). Contrastive decoding necessitates a suitable amateur model for effectiveness [20; 25], but selecting a separate amateur model with same vocabulary is not always feasible. In comparison, SCMoE capitalizes on the MoE models’ inherent strong and weak activation to conduct self-contrast. Different weak activation can be viewed as different amateur models, offering higher flexibility and thus help to find the ideal one for contrast.
@@ 4.3 Combining SCMoE with Self-Consistency
Using self-consistency  [35] for multiple sampling and taking a majority vote to determine the final answer is a common method to improve LLMs’ performance. Therefore, we explore whether SCMoE can combined with self-consistency. For vanilla self-consistency, we use temperature sampling with temperature $\tau=0.7$ to reach the best baseline performance. For self-consistency with SCMoE, we simply employ $\beta=0.5$, rank-3 routing as weak activation, according to the best hyperparameters setting from Table 8. It is worth noting that since SCMoE already has a mask $\alpha=0.1$ to limit the sampling range of the vocabulary, we do not perform any additional temperature processing on the final logits. As shown in Figure 5, SCMoE (67.94) yields comparable results with major@5 (66.87). Furthermore, SCMoE can enhance the major@20 accuracy from 75.59 to 78.31 (+2.72) on GSM8K.
@@ 5 Related Work
The Mixture-of-Experts (MoE) model was initially introduced by A. Jacob et al. [15]. Previous studies have demonstrated that sparsely gated MoE models can significantly improve model capacity and efficiency, enabling superior performance compared to dense ones [30; 10; 39; 40]. In MoE models, a static number of experts are activated regardless of the varying complexity presented by input tokens. Typically, top-1 or top-2 experts are activated in these models [19; 12]. In the era of LLMs, numerous extensive open-source models based on MoE architecture have emerged. Specifically, both Mixtral 8x7B [17] and Grok-1 [37] introduce an 8-expert MoE that uses a top-2 routing algorithm during inference. DeepSeekMoE [9] and QwenMoE [34], on the other hand, both employ a fine-grained expert segmentation, applying 2 shared experts with $N$ routed experts. As a result, they use $k$+2 experts for inference, with 2 fixed shared experts and top-$k$ routed experts.
While several works have attempted to examine pruning or dynamic routing algorithms for MoE models [23; 11; 14] from the perspective of reducing computational costs while maintaining performance. Our approach differs in that we investigate the utilization of unchosen experts in a self-contrast manner to boost MoE models’ capability without increasing too much computation.
@@
The idea of employing contrast to enhance language modeling has been explored through various approaches. Specifically, the contrast enables language models to discern between desirable and undesirable behaviors, a distinction that the conventional maximum log-likelihood modeling often fails to adequately capture [1]. One line of research focuses on training-time optimization. Reinforcement learning from human feedback (RLHF) [32; 4; 27] trains reward models by contrasting the rewards associated with desirable outputs to those of undesirable ones, and then optimize the LLM to maximize rewards through reinforcement learning. RRHF [38], DPO [28], and PRO [31] eliminate the necessity of constructing reward models and instead directly optimize LLMs by contrasting preferred responses versus dispreferred ones. Another research avenue focuses on inference-time optimization. DExperts [22] fine-tunes two models with desirable and undesirable attributes separately, guiding the base model by leveraging the contrast between those models. Contrastive Decoding [20; 25] contrasts base model with an amateur model to mitigate undesirable tendencies of the amateur. Emulated fine-tuning [24] and proxy-tuning [21] achieve training-free alignment in a similar way, applying the contrast between aligned and unaligned models as a reward signal to guide the decoding process of a larger unaligned LLM. Contrastive Search [33] uses a look-ahead contrastive mechanism and penalizes tokens compromising the isotropy of the model’s latent space. DoLa [7] obtains the next-token distribution by contrasting the logits differences between the last layer and a premature layer to improve factuality.
@@ Appendix D Detailed Results of Different Hyperparamters Setting for Each Method
Here, we report the performance of each decoding method in Tables 194, Figure 4 under various hyperparameter settings in 78910.

### v2

@@ 1 Introduction
Scaling up model parameters, dataset size and training time has been considered the most direct and effective approach to improving foundation models’ performance [1, 2, 3]. However, scaling dense models substantially increases computational costs, which poses a significant practical challenge. Mixture-of-Experts (MoE) [4, 5, 6, 7, 8, 9] has emerged as a compelling solution for optimizing the balance between model capacity and computation overhead in the era of large foundation models.
MoE models achieve the goal by sparsely activating only a portion of the parameters for each specific input. Specifically, in MoE models, parameters are grouped into a bunch of experts, MoE models only activate some of them for processing a given input. This selective activation is achieved through a routing mechanism that dispatches each input token to a fixed number of experts (e.g, top-$k$ routing [6, 8, 10, 11]). Therefore, compared to their dense counterparts, MoE models enjoy more efficient training with significantly reduced computational costs [5, 6, 7, 8, 9, 11]. At the inference stage, they typically adhere to the same routing strategy as the training stage, activating only a small fraction of experts. Basically, for each input token, most of the well-trained experts do not contribute to the output prediction. As a result, the potential of utilizing more experts during the inference stage to enhance performance remains underexplored.
@@
In this paper, we investigate the impact of unchosen experts111 Unchosen experts refer to the experts not selected by default routing (e.g., top-2 routing in Mixtral 8x7B). on the performance of MoE models and explore their suitable usage. A direct hypothesis is that incorporating more experts improves MoE models and helps solve more difficult problems [12, 13, 14]. However, in our exploratory experiment on Mixtral 8x7B [6], we find simply raising the number of activated experts (blue lines in Figure 1) does not lead to stable improvements and may even hurt performance on different tasks. This indicates that unchosen experts may contribute little or even negatively to the final performance, which is contrary to the common perception of unchosen experts as candidates of positive power.
@@ 2.2 Divergence Between Different Routing Strategies: An Exploratory Analysis
To sum up, it is observed that, in scenarios demanding reasoning capability for next-token prediction, MoE models with top-2 and rank-$k$ routing strategies demonstrate distinct generation behaviors. This phenomenon suggests that different experts do not always act synergistically, and could in fact exhibit conflicting behaviors. To harness such information introduced by more experts, a feasible approach is to apply contrastive methods [17, 18] to transform the observed negative impacts into positive ones.
@@ 3.3 Results
SCMoE enhances mathematical reasoning by a +5.10 increase on GSM8K, commonsense reasoning by a +3.46 improvement on StrategyQA. Moreover, in code generation, SCMoE gets improvements of +2.60 and +7.92 on the MBPP and HumanEval, respectively. In contrast, traditional search-based methods do not demonstrate substantial improvements on MoE models. In particular, DoLa’s performance not only fails to surpass, but actually falls below the greedy baseline, particularly due to its inability to terminate generation sequences appropriately (for specific examples, refer to Table 11 in the appendix). Meanwhile, contrastive decoding with Mistral 7B as the amateur model does not result in consistent improvements, and even a decrease in pass@1 accuracy on MBPP (-1.00). Contrastive decoding necessitates a suitable amateur model for effectiveness [18, 25], but selecting a separate amateur model with same vocabulary is not always feasible. In comparison, SCMoE capitalizes on the MoE models’ inherent strong and weak activation to conduct self-contrast. Different weak activation can be viewed as different amateur models, offering higher flexibility and thus help to find the ideal one for contrast.
@@ 4.3 Combining SCMoE with Self-Consistency
Using self-consistency  [26] for multiple sampling and taking a majority vote to determine the final answer is a common method to improve LLMs’ performance. Therefore, we explore whether SCMoE can combined with self-consistency. For vanilla self-consistency, we use temperature sampling with temperature $\tau=0.7$ to reach the best baseline performance [27]. For self-consistency with SCMoE, we simply employ $\beta=0.5$, rank-3 routing as weak activation, according to the best hyperparameters setting from Table 8. It is worth noting that since SCMoE already has a mask $\alpha=0.1$ to limit the sampling range of the vocabulary, we do not perform any additional temperature processing on the final logits. As shown in Figure 5, SCMoE (67.94) yields comparable results with major@5 (66.87). Furthermore, SCMoE can enhance the major@20 accuracy from 75.59 to 78.31 (+2.72) on GSM8K.
@@ 5 Related Work
The Mixture-of-Experts (MoE) model was initially introduced by A. Jacob et al. [29]. Previous studies have demonstrated that sparsely gated MoE models can significantly improve model capacity and efficiency, enabling superior performance compared to dense ones [4, 5, 11, 30]. In MoE models, a static number of experts are activated regardless of the varying complexity presented by input tokens. Typically, top-1 or top-2 experts are activated in these models [15, 10]. In the era of LLMs, numerous extensive open-source models based on MoE architecture have emerged. Specifically, both Mixtral 8x7B [6] and Grok-1 [8] introduce an 8-expert MoE that uses a top-2 routing algorithm during inference. DeepSeekMoE [7] and QwenMoE [9], on the other hand, both employ a fine-grained expert segmentation, applying 2 shared experts with $N$ routed experts. As a result, they use $k$+2 experts for inference, with 2 fixed shared experts and top-$k$ routed experts.
While several works have attempted to examine pruning or dynamic routing algorithms for MoE models [31, 32, 14] from the perspective of reducing computational costs while maintaining performance. Our approach differs in that we investigate the utilization of unchosen experts in a self-contrast manner to boost MoE models’ capability without increasing too much computation.
@@
The idea of employing contrast to enhance language modeling has been explored through various approaches. Specifically, the contrast enables language models to discern between desirable and undesirable behaviors, a distinction that the conventional maximum log-likelihood modeling often fails to adequately capture [33]. One line of research focuses on training-time optimization. Reinforcement learning from human feedback (RLHF) [34, 35, 36] trains reward models by contrasting the rewards associated with desirable outputs to those of undesirable ones, and then optimize the LLM to maximize rewards through reinforcement learning. RRHF [37], DPO [38], and PRO [39] eliminate the necessity of constructing reward models and instead directly optimize LLMs by contrasting preferred responses versus dispreferred ones. Another research avenue focuses on inference-time optimization. DExperts [17] fine-tunes two models with desirable and undesirable attributes separately, guiding the base model by leveraging the contrast between those models. Contrastive Decoding [18, 25] contrasts base model with an amateur model to mitigate undesirable tendencies of the amateur. Emulated fine-tuning [40] and proxy-tuning [41] achieve training-free alignment in a similar way, applying the contrast between aligned and unaligned models as a reward signal to guide the decoding process of a larger unaligned LLM. Contrastive Search [23] uses a look-ahead contrastive mechanism and penalizes tokens compromising the isotropy of the model’s latent space. DoLa [24] obtains the next-token distribution by contrasting the logits differences between the last layer and a premature layer to improve factuality.
@@ 6 Conclusion
Acknowledgements
The authors would like to thank Zicheng Lin, Xinzhe Ni, Yifan Wang, and Qingyan Guo for their valuable feedback and discussions. This work was partly supported by the Shenzhen Science and Technology Program (JCYJ20220818101014030) and the "Graph Neural Network Project" of Ping AnTechnology (Shenzhen) Co., Ltd.
@@ Appendix D Detailed Results of Different Hyperparamters Setting for Each Method
There is one fixed value for the hyperparameter $\alpha=0.1$ in Equation 6 that generalizes across various domains. To provide some clarity, when $\alpha$ is set closer to 1, the contrastive process activates fewer vocabulary for strong activation, resulting in minimal changes after the self-contrast. Conversely, setting $\alpha$ closer to 0 allows more vocabulary tokens to be considered in the self-contrast process, leading to significant changes and potentially introducing more noisy information. A suitable $\alpha$ should strike a balance between including ideal tokens, which can lead to accurate results in the contrastive vocabulary, and avoiding the introduction of excessive noise from an overly large vocabulary. Previous work [18] on masking vocabulary based on $\alpha$ suggests that $\alpha=0.1$ is quite robust and generalizes well across various domains. This guides our choice in this setting.
Moreover, we report the performance of each decoding method in Tables 194, Figure 4 under method-specific hyperparameter settings in 78910.
@@ Appendix E Failed Examples of DoLa on Mixtral 7x8B
Appendix F Scope of SCMoE’s Effectiveness
The strength of SCMoE lies in its ability to handle tasks requiring intricate reasoning processes by leveraging both strong and weak activations, which benefits in scenarios demanding reasoning capability for next-token prediction. In contrast, benchmarks like MMLU [42] do not have explicit (verbalized) reasoning paths, which SCMoE is dedicated to helping. Therefore, SCMoE, similar to other generation decoding strategies like contrastive search [23] and contrastive decoding [18], may not exhibit distinct advantages.